Electrophoresis is a medical therapy technique used to deliver medicaments into the patients body with help of an electric current. During the procedure, two oppositely charged plates are applied to the body providing flow of the current through the tissue. Suppose that current used in first five procedures was 7 mA (duration of the procedure 6 min) and during the last fifteen procedures 14 mA (duration 15 min). Find charge that passes through the body during 20 electrophoresis procedures.


1
Expert's answer
2022-02-02T12:13:28-0500

Answer

For I1=7 mA

(duration of the procedure

t1= 6 min=360sec) and

during the last fifteen procedures

I2= 14 mA

(duration t2= 15 min=900sec).

Total charge

"Q=Q_1+Q_2\\=5\times7\times10^{-3}\times360+15\times14\times10^{-3}\times900\\=12.6+189\\=201.6C"